OpenStudy (anonymous):

PLEASE PLEASE PLEASE HELP ME?! Part 1: Explain which companies would advertise APR over APY and describe how this would help them attract customers. (4 points) Part 2: Create a unique APR (state how often the rate is compounded) and calculate the corresponding APY. Use a comparison of the two rates to verify your answer to part 1. (4 points) (8 points)

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

because the apy is larger than the apr

OpenStudy (anonymous):

lets do an example pick an interest rate as a precent

OpenStudy (anonymous):

you want me to pick one?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

ok i pick 6%

OpenStudy (anonymous):

i mean APR is 6%, i.e. the "annual percentage rate" is 6% if the interest is compounded only once a year, then you get an additions 6% of your deposit. say you deposit $100 then you get $6 for a total of $106 but the APY is different

OpenStudy (anonymous):

if the annual percentage rate is 6% but it is compounded more frequently, then you get more money so for example lets compound it monthly and invest the same $100 then in one year you would have \[100(1+\frac{.06}{12})^{12}=100(1.005)^{12}=106.17\]

OpenStudy (anonymous):

not much different, but a little larger so if the APR is 6% then the APY is 6.17%

OpenStudy (anonymous):

without using the $100 you can calculate the APY that example by computing \[(1+\frac{.06}{12})^{12}-1\] and getting \(.0617\) which is \(6.17\%\)

OpenStudy (anonymous):

how would they attract customers though?

OpenStudy (anonymous):

because \(6.17\%\) is larger than \(6\%\)